OpenStudy (anonymous):

round 24.617 to the nearest hundreth

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

$$\Huge 24.617=24+\frac{6}{10}+\frac{1}{100}+\frac{7}{1000}$$

OpenStudy (anonymous):

24.617 nearest hunderdth = 2 numbers after decimal place soo 24.62